OpenStudy (anonymous):

Find the equation of the circle whose center and radius are given. center (0, 8), radius = 8

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@wio ?

OpenStudy (anonymous):

\[ (x-h)^2+(y-k)^2=r^2 \]Use this equation. \((h,k) = (0, 8)\) and \(r=8\)

OpenStudy (anonymous):

(x-0)^2+(y-8)^2=8^2?

OpenStudy (anonymous):

yes but don't wriet \((x-0)^2\) or your teacher will think you are a moron just write \(x^2\)

OpenStudy (anonymous):

okay...Thanks!